OpenStudy (anonymous):

use the associative property of multiplication to write an equivalent expression (cd)e

OpenStudy (anonymous):

http://www.321know.com/pro74bx2.htm

OpenStudy (anonymous):

nice thanks

OpenStudy (anonymous):

associative: (ab)c=a(bc) so for this, (cd)e=c(de)

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

anytym ! :-)

OpenStudy (anonymous):

(cd)e=c(de) anjali is right

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Anjali is always right! ECO(H + DU :D

OpenStudy (anonymous):

dont mention its wht i do!